# Decentralizing Test-time Adaptation under Heterogeneous Data Streams

### v1

@@ 4.2.1 Frequency-based Decentralized learning
Frequency Feature Extraction. We start by extracting frequency domain features from the input images to identify distinct distribution shifts. Let denote a batch of input images, where is the batch size, is the number of channels, and are the height and width of the images. We first apply a Fourier transform to each image to obtain its frequency domain representation . Particularly, we focus on the amplitude spectrum in Eq. 1, filtering out low-frequency elements using mask to emphasize the high-frequency components that are more likely to indicate shifts in distribution:
@@
Frequency-Based Clustering. We then employ a clustering algorithm (e.g., K-means) to partition the frequency features into clusters, each corresponding to a different type of distribution shift. The process is formalized as:
@@
where , , denotes the 1D high-frequency component of the amplitude spectrum (the centroids of the clusters and the cluster assignments for each image), refers to high-frequency components, and is flattened dimension.
@@
Decentralized Fine-tuning. Test-time fine-tuning is then decentralized across these clusters, allowing for specialized adaptation within each subgroup: For each cluster , we adapt a specialized model that is fine-tuned using only the data within that cluster:
@@
where represents the data distribution within cluster , and is the loss function.
@@
Parameter Aggregation. To integrate knowledge from all subnetworks and prevent degradation on specific subdomains, we perform an aggregation of their parameters:
@@
where denotes the number of samples in cluster . This aggregation step combines the parameter updates from each subnetwork proportionally to its cluster size. The updated global parameters are then distributed back to each subnetwork, updating its parameters as follows:
4.2.2 Frequency-based Augmentation
Insight: Although decentralized learning effectively handles data heterogeneity within the current batch, it may still suffer from inadequate characterization of each distribution shift due to limited batch data, specifically considering the uniform amplitude distributions resulting from prior clustering. Typically, TTA methods attempt to enhance the overall quality of observed target samples via data augmentation. However, traditional augmentations in TTA, borrowed from standard computer vision practices such as rotation, clipping, and mixup, albeit beneficial in scenarios with single distribution shifts, struggle to guarantee targeted fine-tuning under more complex, mixed distribution shift scenarios.
Solution: To overcome these limitations, we propose a frequency-based augmentation strategy tailored for TTA under mixed distribution shifts. Our method specifically perturbs the amplitude components of each sample in Fourier space. This targeted approach allows us to augment samples comprehensively, enhancing data quality for each individual shifting case and boosting mode performance across complex real-world scenarios.
@@ 4.2.2 Frequency-based Augmentation
Sample Selection Mechanism. We leverage a criterion derived from the weighted entropy framework used in ETA [26] based on two primary conditions:
@@
Frequency-Based Augmentation. The augmentation process involves perturbing the amplitude spectrum. Let represent the amplitude spectrum of a selected sample . To generate a perturbed amplitude spectrum , we apply a random Gaussian perturbation:
@@
Loss Function. The training objective combines the entropy loss of the selected samples with a consistency loss from the augmented samples. The total loss is defined as:
where the entropy loss for the original samples is given by with being the predicted probability over the classes for the sample , and the consistency loss is defined as the cross-entropy between the prediction of the augmented sample and the pseudo-label from the original sample: .
@@
| Methods | C10 | C100 | IN(BN) | IN(LN) |
| Source | 43.5 | 46.5 | 82.0 | 60.2 |
| TBN | 79.2 | 92.3 | 94.2 | - |
| TENT (ICLR 21’) | 86.6 | 98.4 | 99.5 | 77.9 |
| ETA (ICML 22’) | 86.1 | 96.2 | 99.7 | 73.9 |
| AdaContrast (CVPR 22’) | 69.8 | 73.2 | 98.5 | 94.9 |
| CoTTA (CVPR 22’) | 82.7 | 92.8 | 98.0 | 92.6 |
| SAR (ICLR 23’) | 78.8 | 95.8 | 98.2 | 54.0 |
| RoTTA (CVPR 23’) | 64.6 | 65.3 | 89.3 | 74.2 |
| RDumb (NeurIPS 23’) | 86.2 | 98.4 | 98.1 | 56.5 |
| DeYO (ICLR 24’) | 87.0 | 98.1 | 99.1 | 52.0 |
| UnMix-TNS (ICLR 24’) | 41.9 | 50.1 | 84.3 | - |
| FreDA (ours) | 23.0 | 34.7 | 67.2 | 48.7 |
5 Experiments
Datasets and Experimental Settings. To provide a comprehensive evaluation of TTA deployment, we test models over multiple datasets under three different scenarios:
Natural Domain Shifts: We extend our evaluation to DomainNet126 [29], which presents natural shifts across four domains (Real, Clipart, Painting, Sketch) encompassing 126 classes, representing a subset of the larger DomainNet dataset.
For corruption datasets, the model is pretrained on the clean dataset and the 15 corruptions are randomly mixed as the target distribution. We leverage the highest severity = 5 in all the experiments. In DomainNet126 and Camelyon17, one subdomain is selected as the source, and the others serve as mixed target distributions. More implementation details are provided in Appendix A. All reported results are averaged over runs with fixed seeds (0, 1, and 2).
| DomainNet126 Methods Real Painting Clipart Sketch Avg. Source 45.2 41.6 49.5 45.3 45.4 TBN 45.5 39.9 45.9 37.5 42.2 TENT (ICLR 21’) 42.2 37.8 44.7 37.5 40.6 ETA (ICML 22’) 41.1 37.3 43.4 36.4 39.5 SAR (ICLR 23’) 43.2 38.5 44.8 37.0 40.9 DeYO (ICLR 24’) 40.9 36.4 43.6 36.9 39.4 FreDA (ours) 40.2 36.1 40.0 33.6 37.5 | Camelyon17 C1 C2 C3 C4 C5 Avg. 21.6 43.6 52.5 47.4 47.6 42.5 26.5 38.5 31.7 39.4 32.8 33.8 44.7 50.5 49.9 49.1 48.6 48.6 47.4 52.5 47.9 49.9 39.2 47.4 26.5 38.5 31.7 39.4 32.8 33.8 50.4 50.3 48.8 51.7 50.5 50.4 18.6 24.7 24.8 40.5 30.8 27.9 |
Adaptation Scenarios. To evaluate models in adapting to heterogeneous data streams, we focus on two primary distribution shift scenarios including:
Mixed Domains: A long test sequence where consecutive test samples may come from different domains.
Mixed Domains & Dependent Sampling: Extends the mixed distribution framework by introducing sequential, time-correlated data from the same class.
The detailed description of the settings can be found in Appendix E. Moreover, While our primary focus is on mixed domains, we have also included the commonly used continual setting for evaluation in Appendix C.
Baselines. We compare FreDA with 10 models: TBN [25], TENT [34], CoTTA [36], ETA [26], SAR [27], AdaContrast [2], RoTTA [45], RDumb [28], DeYO [17], and UnMix-TNS [32]. See more information in Appendix B.
FreDA consistently improves across different distribution shifts. Our method consistently attains the lowest classification error rates across all evaluated datasets (see Table 1 and 3). Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%. This significant improvement is particularly notable where other approaches falter if compared with no training (TBN), meaning they struggle to adapt to the complex medical imaging data. By effectively handling high variability and intricate patterns in the data, FreDA maintains superior accuracy and adaptability.
FreDA effectively mitigates both covariate and label shifts. In environments characterized by simultaneous covariate and label shifts, our approach keep showing exceptional adaptability (see Table 2). We attribute this success to FreDA’s ability to separate covariate shifts from label shifts via decentralized learning. FreDA achieves this by first isolating target different distribution shifts and then focus on learning label shifts for each specific distribution. This sequential approach prevents models from being overwhelmed by simultaneous shifts, allowing it to address each type of shift independently and effectively.
FreDA remains stable under various batch size. To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distribution shifts. In Figure 3, we present the results on CIFAR-10-C, CIFAR-100-C, and ImageNet-C datasets using batch sizes ranging from 200 (64) down to 1. Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance. This stability under limited batch sizes demonstrates FreDA’s robustness, making it highly suitable for real-world applications where large batches is not always feasible.
| DT | SS | SA | C10 | C100 | IN(BN) | IN(LN) |
| 44.1 | 82.5 | 86.4 | 55.0 | |||
| ✓ | 24.8 | 54.2 | 81.2 | 95.2 | ||
| ✓ | 29.6 | 37.5 | 71.0 | 51.1 | ||
| ✓ | 39.4 | 71.7 | 92.9 | 59.5 | ||
| ✓ | ✓ | 24.3 | 36.3 | 69.4 | 49.6 | |
| ✓ | ✓ | 27.7 | 36.2 | 65.9 | 50.1 | |
| ✓ | ✓ | 24.4 | 50.2 | 77.7 | 95.3 | |
| ✓ | ✓ | ✓ | 22.9 | 34.7 | 67.2 | 48.8 |
FreDA enhances adaptation via synergistic designs. This section validates our designs by ablating its three key modules – Decentralized Training (DT), Sample Selection (SS), and Sample Augmentation (SA). The baseline here leverages only the entropy loss. From Table 4, we have the following observations: 1) Implementing decentralized training alone results in substantial improvements, reducing error rates dramatically across all datasets. 2) The impact of sample selection varies across datasets. While significantly improving performance on CIFAR100-C, it increase error rate on Camelyon. This variation suggests that sample selection helps the model focus on more representative or challenging samples but may not be effective across all datasets, highlighting its dataset-specific nature. 3) Sample augmentation alone tends to increase error rates, suggesting that although this approach introduces useful variability, it may introduce unexpected noise under the absence of proper selection or decentralized training. 4) The combined approach delivers the best performance across all datasets, showing the synergistic effect of our different designs.
6 Conclusion
This paper advances Test-Time Adaptation (TTA) by tackling the real-world complexities of heterogeneous data streams. Our decentralized approach precisely manages diverse data shifts, improving model adaptation in varied settings. By integrating Fourier-based augmentation, we expand the range of confident samples for each distribution shift, further boosting model performance. The experimental results underscore the efficacy of FreDA, highlighting its potential to influence the field and guide future research in adapting to dynamic and diverse data shifts.
Supplementary Material
A Implementation Details
Pretrained Models. We utilize models from RobustBench [3], including WildResNet-28 [46] for CIFAR-10-C and ResNeXt-29 [37] for CIFAR-100-C, both pretrained by Hendrycks et al. [12]. For ImageNet-C, the pretrained ResNet-50 [10] and VitBase-LN [5] are obtained from torchvision. For DomainNet126, pretrained ResNet-50 are sourced from AdaContrast [2], while for Camelyon17, we train a DenseNet-121 [14] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [16].
Hyperparameter Configuration. The batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods. The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively. The learning rate is proportionally decreased in the experiment studying the effect of batch size. The Kmeans Size is 512, Clutser Number is 4, Communication Interval is 10 across all the tasks. The perutrbation magnitude is fixed to 0.1 and the coefficient in loss function is fixed to 0.5. The parameter controlling the dependent sampling (Dirichlet distribution) is set to 0.1 for CIFAR10-C and adjusted to 0.01 for CIFAR100-C, ImageNet-C following UnMix-TNS [32]. Two threshold in Eq. 7 is set to the same value for corruption datasets and DomainNet126 following ETA [26]. While for Camelyon17, the class diversity related threshold is adjusted to 0.9 empirically.
B Compared Methods
TBN [25] re-estimates batch normalization statistics from test data. TENT [34] minimizes prediction entropy to optimize batch normalization. CoTTA [36] addresses long-term test-time adaptation in changing environments. ETA [26] and SAR [27] exclude unreliable and redundant samples during optimization. AdaContrast [2] utilizes contrastive learning to refine pseudo-labels and improve feature learning. RoTTA [45] presents a robust batch normalization scheme with a memory bank for category-balanced estimation. RDumb [28] leverages weighted entropy and periodically resets the model to its pretrained state to prevent collapse. DeYO [17] quantifies the impact of object-destructive transformations for sample selection and weighting. UnMix-TNS [32] introduces a test-time normalization layer for non-i.i.d. environments by decomposing BN statistics. For fair comparisons, we conduct experiments using the open source online TTA repository [4]111https://github.com/mariodoebler/test-time-adaptation, which provides codes and configurations of state-of-the-art TTA methods.
C Continual Setting Evaluation
| Methods | C10 | C100 | IN(BN) | IN(LN) |

### v2

@@ V-A1 Frequency Feature Extraction
We start by extracting frequency domain features from the input images. Let denote a batch of input images, where is the batch size, is the number of channels, and are the height and width. We first apply a Fourier transform to each image . This transform converts the image from the spatial domain to the frequency domain, producing a complex-valued representation that contains both real and imaginary part . Next, we compute the amplitude spectrum using . It reveals the intensity of the frequency content, e.g., high-frequency amplitudes highlight edges and fine details while low-frequency amplitudes emphasize the overall structure and gradual changes. Then, we filter out low-frequency elements using mask to emphasize the high-frequency components that are more likely to indicate shifts in distribution:
V-A2 Frequency-Based Clustering
@@
We then employ a clustering algorithm (e.g., K-means) to partition the frequency features into clusters, each corresponding to a different type of distribution shift. The process is formalized as:
@@
where , , denotes the 1D high-frequency component of the amplitude spectrum, the centroids of the clusters and the cluster assignments for each image. refers to high-frequency components, and is flattened dimension.
V-A3 Decentralized Fine-tuning
@@
Test-time fine-tuning is then decentralized across these clusters, allowing for specialized adaptation within each subgroup: For each cluster , we adapt a specialized model that is fine-tuned using only the data within that cluster:
@@
where represents the data distribution within cluster , and is the loss function. The predictions of each iteration are collected and sorted after the local fine-tuning.
V-A4 Base Model Enhancement
@@
To integrate knowledge from all subnetworks and prevent degradation on specific subdomains, we periodically aggregation their parameters at intervals of time :
@@
where denotes sample number in cluster . This aggregation step combines the parameter updates from each subnetwork proportionally to its cluster size. The updated parameters are then distributed back to each subnetwork, initializing them for the next batch of training: .
V-B Frequency-based Augmentation
While decentralized learning effectively mitigates batch-level heterogeneity, it may falls short in accurately characterizing individual distribution shifts – largely due to limited batch size and noise introduced by coarse clustering. To improve target data quality, TTA methods commonly adopt data augmentation as a practical strategy to enhance model generalization. However, conventional augmentation techniques – typically borrowed from standard computer vision tasks such as rotation, cropping, and mixup – are primarily tailored for single-shift scenarios. They often fail to provide the targeted, distribution shifts-aware augmentation as required under mixed domain shifts. To this end, we propose a frequency-based augmentation strategy tailored for TTA in mixed-shift environments. Specifically, our method perturbs the amplitude components of target samples in the Fourier space, enabling targeted augmentation with respect to each sample’s underlying distribution. By manipulating the frequency-domain characteristics, our approach generates diverse yet distribution-consistent variations, thereby enriching subdomain representations and facilitating robust model adaptation
V-B1 Sample Selection Mechanism
@@
We first select the reliable samples in each local model leveraging a criterion derived from the weighted entropy framework used in ETA [6] based on two primary conditions:
@@
V-B2 Frequency-Based Augmentation
The augmentation process involves perturbing the amplitude spectrum. Let represent the amplitude spectrum of a selected sample . To generate a perturbed amplitude spectrum , we apply a random Gaussian perturbation:
@@ V-B2 Frequency-Based Augmentation
V-B3 Loss Function
The training objective combines the entropy loss of the selected samples with a consistency loss from the augmented samples. The total loss is defined as:
where the entropy loss for the original sample is given by with being the predicted probability over the classes, and the consistency loss is defined as the cross-entropy between the prediction of the augmented sample and the pseudo-label from the original sample.
V-C Overall Framework
We provide the overall pipeline algorithm of FreDA in Algorithm 1. During implementation, we leverage a memory bank strategy [25, 4, 32, 33] that is updated in real time. This design serves two purposes: 1) to ensure accurate clustering—since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
VI Theoretical Insights
To facilitate theoretical understanding, we leverage the expansion-based analysis framework proposed in [34]. We first revisit the core concepts and then highlight the theoretical advantages of our method under mixed domain shifts.
VI-A Expansion Theory
Definition 1 (-expansion).
A class-conditional distribution satisfies -expansion if with :
where refers to the neighborhood of under data augmentations.
Definition 2 (Separation).
is -separated if:
where is an -ball of radius .
The core theorem from [34] states:
Theorem 1 (Pseudo-label Denoising).
Under -expansion and -separation, any classifier of:
achieves error:
where is pseudo-label loss and is consistency regularizer.
VI-B Analysis of Our Method
Our method reduces both terms in Theorem 1 to achieve a tighter bound through two mechanisms:
VI-B1 Frequency-Coherent Partitioning.
Let be the sub-domains identified by frequency-based partitioning. For measurable sets where (the support of ), define:
VI-B2 Augmentation-Induced Expansion
Our frequency augmentation expands neighborhoods:
where is the base augmentation neighborhood. It is worth noting that, unlike most methods, we do not apply base augmentation. The term is retained in the formulation for theoretical completeness. Our choice of this representation is intended to highlight the expansion gain brought by incorporating frequency augmentation.
The expansion gain becomes:
Substituting into Theorem 1, the error becomes:
Notation Summary
: Expansion factor (Def. 1).
Class-wise connectivity metric: Larger implies stronger neighborhood propagation of local consistency to global predictions.
: Inter-class separation probability (Def. 2).
Robustness measure: Probability of different classes having overlapping neighborhoods under perturbations, lower indicates clearer decision boundaries.
: Consistency regularizer (Thm. 1).
Stability term: Penalizes prediction inconsistency between original inputs and their augmented variants.
: Error reduction in -th sub-domain.
Specialization gain: Reduced pseudo-label noise in sub-domain due to frequency-based partitioning.
: Expansion gain from frequency augmentation.
Neighborhood enhancement: Additional expansion capability measured as the relative increase of augmented neighborhoods.
@@ VI-B2 Augmentation-Induced Expansion
VII Experiments
VII-A Datasets and Experimental Setup
VII-A1 Datasets
To provide a comprehensive evaluation of TTA deployment, we test models over multiple datasets under three different scenarios:
Natural Domain Shifts: We extend evaluation to DomainNet126 [35], which presents natural shifts across four domains (Real, Clipart, Painting, Sketch) encompassing 126 classes as a subset of the larger DomainNet dataset.
For corruption datasets, the model is pretrained on the clean dataset and the 15 corruptions are randomly mixed as the target distribution. We leverage the highest severity = 5 in all the experiments. In DomainNet126 and Camelyon17, one subdomain is selected as the source, and the others serve as mixed target distributions. All reported results are averaged over runs with fixed seeds (0, 1, and 2).
VII-A2 Baselines
We compare FreDA with 10 models, including TBN [31], TENT [3], CoTTA [5], ETA [6], SAR [7], AdaContrast [4], RoTTA [25], RDumb [9], DeYO [10], and UnMix-TNS [28]. TBN [31] re-estimates batch normalization statistics from test data. TENT [3] minimizes prediction entropy to optimize batch normalization. CoTTA [5] addresses long-term test-time adaptation in changing environments. ETA [6] and SAR [7] exclude unreliable and redundant samples during optimization. AdaContrast [4] utilizes contrastive learning to refine pseudo-labels and improve feature learning. RoTTA [25] presents a robust batch normalization scheme with a memory bank for category-balanced estimation. RDumb [9] leverages weighted entropy and periodically resets the model to its pretrained state to prevent collapse. DeYO [10] quantifies the impact of object-destructive transformations for sample selection and weighting. UnMix-TNS [28] introduces a test-time normalization layer for non-i.i.d. environments by decomposing BN statistics. For fair comparisons, we conduct experiments using the open source online TTA repository [37]111https://github.com/mariodoebler/test-time-adaptation, which provides codes and configurations of state-of-the-art TTA methods.
VII-A3 Pretrained Models
We utilize models from RobustBench [38], including WildResNet-28 [39] for CIFAR-10-C and ResNeXt-29 [40] for CIFAR-100-C, both pretrained by [41]. For ImageNet-C, the pretrained ResNet-50 [42] is obtained from torchvision. For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [43] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [44].
VII-A4 Hyperparameter Configuration
The batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods. The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively. The learning rate is proportionally decreased in the experiment studying the effect of batch size. The Kmeans Size is 512, Clutser Number is 4, Communication Interval is 10 across all the tasks. The perturbation magnitude is fixed to 0.1 and the coefficient in loss function is fixed to 0.5. Two threshold in Eq. 5 is set to the same value for corruption datasets and DomainNet126 following ETA [6]. While for Camelyon17, the class diversity related threshold is adjusted to 0.9 empirically.
VII-B Main Results
VII-B1 FreDA improves across diverse distribution shifts
Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE V and III). Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%. This significant improvement is particularly noTABLE where other approaches falter if compared with no training (TBN), meaning they struggle to adapt to the complex medical imaging data. By effectively handling high variability and intricate patterns in the data, FreDA maintains superior accuracy and adaptability.
VII-B2 FreDA remains stable under various batch size
To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distributions. In Figure 5, we present the results on CIFAR-10-C, CIFAR-100-C, and ImageNet-C using batch sizes ranging from 200 (64) down to 1. Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance. This stability demonstrates FreDA’s robustness, making it highly suiTABLE for real-world applications where large batches is not always feasible.
| DT | SS | SA | C10 | C100 | IN |
| 44.1 | 82.5 | 86.4 | |||
| ✓ | 24.8 | 54.2 | 81.2 | ||
| ✓ | 29.6 | 37.5 | 71.0 | ||
| ✓ | 39.4 | 71.7 | 92.9 | ||
| ✓ | ✓ | 24.3 | 36.3 | 69.4 | |
| ✓ | ✓ | 27.7 | 36.2 | 65.9 | |
| ✓ | ✓ | 24.4 | 50.2 | 77.7 | |
| ✓ | ✓ | ✓ | 22.9 | 34.7 | 67.2 |
VII-B3 FreDA enhances adaptation via synergistic designs
This section validates our designs by ablating three key modules – Decentralized Training (DT), Sample Selection (SS), and Sample Augmentation (SA). The baseline here leverages only the entropy loss. From TABLE VI, we have the following observations: 1) Implementing decentralized training and sample selection alone results in substantial improvements, reducing error rates dramatically across all datasets. 2) Sample augmentation alone has the possibility to increase error rates, suggesting that although this approach introduces useful variability, it may introduce unexpected noise under the absence of proper selection or decentralized training. 3) The combined approach delivers the best performance across all datasets, showing the synergistic effect of our different designs.
@@ VII-B3 FreDA enhances adaptation via synergistic designs
VII-C More analysis and Discussion
VII-C1 Performance with Transformer Backbone
In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone. Here, we report results on the ImageNet-C benchmark using ViTBase-LN [45] (see TABLE V), where the pretrained weights are obtained from torchvision. Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments. As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
| CLUSTER_NUM | 2 | 4 | 8 | 16 |
| CIFAR10-C | 23.0 | 22.9 | 23.2 | 24.7 |
| CIFAR100-C | 34.8 | 34.7 | 34.7 | 35.6 |
| IN-C (ResNet) | 68.6 | 67.2 | 67.1 | 70.5 |
| IN-C (ViT) | 50.3 | 48.8 | 49.9 | 50.0 |
| KMEANS_SIZE | 256 | 512 | 1024 | 2048 |
| CIFAR10-C | 23.0 | 22.9 | 23.0 | 22.9 |
| CIFAR100-C | 34.6 | 34.7 | 34.8 | 34.8 |
| IN-C (ResNet) | 69.0 | 67.2 | 67.6 | 67.0 |
| IN-C (ViT) | 49.0 | 48.8 | 48.7 | 48.8 |
| COMM_INTERVAL | 1 | 10 | 100 | 1000 |
| CIFAR10-C | 22.6 | 22.9 | 22.6 | 22.0 |
| CIFAR100-C | 34.7 | 34.7 | 34.9 | 43.2 |
| IN-C (ResNet) | 67.1 | 67.2 | 67.2 | 67.4 |
| IN-C (ViT) | 48.4 | 48.8 | 48.8 | 48.7 |
| PERT_MAGNITUDE | 0.0 | 0.1 | 0.2 | 0.3 |
| CIFAR10-C | 24.3 | 22.9 | 22.5 | 22.2 |
| CIFAR100-C | 36.3 | 34.7 | 34.9 | 34.9 |
| IN-C (ResNet) | 69.3 | 67.2 | 67.0 | 66.9 |
| IN-C (ViT) | 49.6 | 48.8 | 48.5 | 48.6 |
VII-C2 Sensitivity Study
We investigate the impact of the key hyperparameters: CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL, and PERT_MAGNITUDE. From TABLE VI, we have the following observations: 1) The choice of CLUSTER_NUM influences model performance, especially on more complex datasets. While performance remains relatively stable even with a small number of clusters (e.g., two clusters), increasing the number of clusters beyond four tends to lead to slight performance degradation, particularly on datasets like ImageNet-C. This suggests that while more clusters can improve adaptation capacity, excessively increasing the number can introduce overfitting and reduce generalization. Therefore, a moderate number of clusters, around 4, appears to strike a good balance between adaptation flexibility and maintaining model robustness. 2) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes. 3) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 1 to 1,000). On simpler datasets such as CIFAR10-C, infrequent communication (e.g., interval ) performs best, likely due to effective independent learning. In contrast, for complex datasets, more frequent communication (e.g., ) improves performance, likely by mitigating divergence among local branches and ensuring model consistency. 4) Adjusting the perturbation magnitude (PERT_MAGNITUDE) has a noticeable effect, particularly on simpler datasets. Increasing perturbation improves performance, suggesting that spectral augmentation contributes to better generalization. However, for more complex datasets such as CIFAR100-C, higher perturbation levels lead to slight decline in performance, likely due to noise disrupting the alignment of finer features. A perturbation magnitude of 0.1 appears to strike an effective balance, offering the benefits of augmentation without negatively impacting feature alignment.
| Methods | C10 | C100 | IN(ResNet) | IN(ViT) |